Suppose that John is offered a simple gamble in which a single die is rolled. If the die comes up 1 or 2, he wins $60, and if it comes up 3,4,5 or 6 he loses $20. John is just indifferent between taking this simple gamble and not taking it.

(a) What is the expected value of this gamble?

(b) Can you tell from this information whether John is risk averse, risk neutral, or risk loving? Explain.

(c) What is the risk premium that John is being paid in order to accept this gamble?
Answer #1

a) Expected value in the gamble = E

The probability of getting 1 or 2 = [2/6] = 1/3

The probability of getting 3,4,5,6 = 4/6 = 2/3

The winning amount = 60$

The loosing amount = -20$

Expected value = (1/3)x60 + (2/3) x -20

Expected value = 6.67$

b) It is risk averse as it is having a dimnishing marginal utility of wealth every time u roll a dice the probability of loosing is more that winning so it is risk aversing.